# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ 4 Experiments
Baselines & Metrics. We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several existing flow matching baselines. As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings. We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors. For evaluation, we employed the 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, as well as MMD and Energy Distance for statistical fidelity. Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset. Exact metric definitions are present in Appendix B.2. To account for randomness, all reported metrics are averaged across $10$ runs.
Table 3: Extrapolation on the last holdout time point for the Brain Regeneration dataset.
@@
CTF-C 11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
Sampling. A trained velocity field can be evaluated through the samples it generates. We consider two variants of sampling configurations: initial value problem sampling (IVP) and next-step sampling (Next Step). In particular, IVP integrates the learned gradients from the first observed batch of cells and evolves them to a later time point. IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps. In contrast, Next Step integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
4.1 Axolotl Brain Regeneration
We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022). The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage. For our CTF-C method, we present the best ablated $\alpha$, with full ablation results across different $\alpha$ values in Appendix H.
For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and next-step sampling. Table 2 presents the results. Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth. CTF-H consistently achieves the best or comparable performance relative to CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values. This highlights the computational efficiency and superior generalization ability of CTF-H, as it avoids the need for additional hyperparameter tuning while maintaining strong performance.
For extrapolation, we evaluate generation at the last holdout time point, which represents the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon. As shown in Table 3, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate. This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon. Figure 4 in Appendix G.1 further illustrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
Table 4: Interpolation (time 5) and extrapolation (time 8) results on the Organogenesis dataset.
Method $\bm{\lambda}$ $\bm{\alpha}$ Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
@@ 4.1 Axolotl Brain Regeneration
CTF-C 11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $3.137\pm 0.407$ $4.093\pm 1.187$ $1.685\pm 0.096$ $1.714\pm 0.160$
0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $2.938\pm 0.476$ $3.904\pm 1.120$ $1.773\pm 0.053$ $1.880\pm 0.180$
$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $3.200\pm 0.403$ $3.555\pm 0.637$ $1.768\pm 0.058$ $1.858\pm 0.120$
11 $0.2$ $1.880\pm 0.020$ $1.922\pm 0.078$ $3.260\pm 0.880$ $5.264\pm 3.060$ $1.683\pm 0.058$ $1.803\pm 0.117$
0 $0.2$ $1.900\pm 0.035$ $1.912\pm 0.057$ $2.953\pm 0.425$ $3.816\pm 0.970$ $1.715\pm 0.123$ $1.860\pm 0.267$
CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$ $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$
11 $1.898\pm 0.029$ $1.866\pm 0.097$ $5.200\pm 0.799$ $6.306\pm 1.037$ $1.890\pm 0.046$ $1.877\pm 0.103$
$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
Table 5: Interpolation results on the middle holdout time point for the Liver Regeneration dataset.
@@
4.2 Mouse Embryo Organogenesis
We further evaluated ContextFlow on the larger Mouse Organogenesis Spatiotemporal Atlas (MOSTA) Stereo-seq dataset (Chen et al., 2022) spanning measurements from $8$ developmental time points. For the interpolation study of this dataset, we held out time point 5 during training and evaluated its generation during testing. Table 4 shows the evaluation results. We observe that ContextFlow, with both integration strategies, outperforms MOTFM across all metrics, showcasing the effectiveness of the contextual information. While CTF-C shows stronger performance under next-step sampling—albeit only after fine-tuning the trade-off parameter $\alpha$—CTF-H consistently outperforms it in the more challenging IVP-Sampling setting. On the extrapolation task, integrating to the final time point, CTF-H again achieves the strongest performance, underscoring that the entropy-regularized formulation not only removes the need for additional parameter tuning but also offers more robust generalization to unseen temporal horizons.
@@ 4.2 Mouse Embryo Organogenesis
Figure 2: Comparisons of predicted and true cell type distributions.
Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM. In both cases, CTF exhibits lower divergence on average across time points, indicating that the trajectories generated by our model better preserve the biological composition of cell types over time. The cell type progression is further visualized in Figure 10 (Appendix H.6). We show that ContextFlow predicts temporal trajectories of cell types that evolve smoothly and consistently across consecutive developmental stages. Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development progresses, while terminal fates, including muscle, the cartilage primordium and the liver, emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout, confirming that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
4.3 Liver Regeneration
Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages. Unlike earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells. Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance. Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide. Table 5 summarizes the comparisons between MOTFM and our methods. Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.

### v3

@@ 5 Experiments
Baselines & Metrics. We evaluate ContextFlow with PACM and PAER (CTF-C and CTF-H) against several state-of-the-art FM baselines, including conditional flow matching (CFM) (Lipman et al., 2023), minibatch-OT flow matching (MOTFM) (Tong et al., 2024), metric flow matching (Metric-FM) (Kapuśniak et al., 2024), and meta flow matching (MFM) (Atanackovic et al., 2024). To evaluate statistical fidelity, we measure $2$-Wasserstein distance ($\mathcal{W}_{2}$), maximum mean discrepancy (MMD), and energy distance (Energy). To assess the biological plausibility of inferred dynamics, we employ a cell-type-weighted $2$-Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset. Formal definitions of these evaluation metrics are detailed in Appendix A. To account for randomness, all reported metrics are averaged across $10$ runs.
Table 1: Comparisons under interpolation at the middle holdout time point on Brain Regeneration.
Sampling Method Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
Metric-FM $2.697\pm 0.044$ $2.688\pm 0.024$ $\bm{0.029\pm 0.001}$ $9.262\pm 0.311$
MFM $5.809\pm 0.140$ $5.851\pm 0.156$ $0.124\pm 0.002$ $49.770\pm 0.479$
CTF-C ($\lambda=0,\alpha=0.2$) $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
CTF-H ($\lambda=1$) $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $0.030\pm 0.004$ $\bm{6.359\pm 1.336}$
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
Metric-FM $4.319\pm 0.299$ $4.501\pm 0.409$ $0.226\pm 0.007$ $34.941\pm 5.063$
MFM $5.976\pm 0.133$ $5.699\pm 0.311$ $0.177\pm 0.010$ $51.024\pm 0.859$
CTF-C ($\lambda=0,\alpha=0.2$) $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
CTF-H ($\lambda=1$) $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
Sampling. A learned velocity field can be evaluated through the samples it generates. We consider two configurations: initial-value-problem (IVP) and next-step (Next Step) sampling. In particular, IVP integrates the gradients learned from the first observed batch of cells and evolves them to a later time point, providing the most comprehensive evaluation of flow quality, as errors can accumulate across steps. Next Step integrates the gradient only from the most recent batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
5.1 Main Results on Axolotl Brain Regeneration
First, we evaluate our ContextFlow method on longitudinal Stereo-seq spatial transcriptomic data from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022). The dataset contains samples from $5$ developmental stages, with replicates collected from different individual organisms at each stage. For our CTF-C method, we present the best ablated $\alpha$, with full ablation results across different $\alpha$ values in Appendix J.
Interpolation. For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and Next-Step sampling. Table 1 presents the results. Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth. CTF-H consistently achieves the best or comparable performance relative to baselines and CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values. This highlights the computational efficiency and superior generalization of CTF-H, which avoids the need for additional hyperparameter tuning while maintaining strong performance.
Table 2: Comparisons under extrapolation on the last holdout time point on Brain Regeneration.
Sampling Method Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
@@ 5.1 Main Results on Axolotl Brain Regeneration
Metric-FM $8.213\pm 4.959$ $6.972\pm 2.557$ $0.124\pm 0.034$ $29.155\pm 13.973$
MFM $13.267\pm 1.802$ $13.513\pm 1.908$ $0.080\pm 0.001$ $109.714\pm 26.407$
CTF-C ($\lambda=0.5,\alpha=0.5$) $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
CTF-H ($\lambda=1$) $\bm{5.277\pm 0.936}$ $\bm{6.021\pm 1.192}$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
Extrapolation. For extrapolation, we evaluate the data generation at the last holdout time point, which represents the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon. As shown in Table 2, CTF-H again achieves the best overall performance, particularly under IVP sampling, where errors are most likely to accumulate. This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon. Figure 5 in Appendix G.1 further illustrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
Table 3: Comparisons under interpolation (time 55) and extrapolation (time 88) on Organogenesis.
Method Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
@@
Metric-FM $1.897\pm 0.036$ $1.864\pm 0.084$ $4.030\pm 1.552$ $3.826\pm 0.767$ $1.595\pm 0.114$ $1.692\pm 0.272$
MFM $\bm{1.839\pm 0.065}$ $1.901\pm 0.099$ $476\pm 0.371 3.3$$319\pm 0.281$ $1.607\pm 0.049$ $1.407\pm 0.072$
CTF-C ($\lambda=1,\alpha=0.5$) $1.865\pm 0.030$ $\bm{1.852\pm 0.093}$ $3.137\pm 0.407$ $4.093\pm 1.187$ $1.685\pm 0.096$ $1.714\pm 0.160$
CTF-H ($\lambda=0.5$) $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
CTF-H ($\lambda=0$) $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$ $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$
Table 4: Comparisons under interpolation on the middle holdout time point on Liver Regeneration.
@@
5.2 Additional Evaluations on Organogenesis and Liver Regeneration
Organogenesis. We further evaluate ContextFlow on the larger Mouse Organogenesis Spatiotemporal Atlas (MOSTA) Stereo-seq dataset (Chen et al., 2022) spanning measurements from $8$ developmental time points. For the interpolation study of this dataset, we hold out the time point $5$ during training and evaluate its generation during testing. Table 3 reports the evaluation results. We observe that ContextFlow, with both integration strategies, performs best overall relative to baselines across all metrics, showcasing the effectiveness of the contextual information. While CTF-C shows stronger performance under next-step sampling—albeit only after fine-tuning the trade-off parameter $\alpha$—CTF-H consistently outperforms it in the more challenging IVP sampling setting. On the extrapolation task, integrating to the final time point, CTF-H again achieves the strongest performance, underscoring that the entropy-regularized formulation not only removes the need for additional parameter tuning but also offers more robust generalization to unseen temporal horizons.
Liver Regeneration. Lastly, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), which is collected across $3$ distinct regeneration stages. Unlike previous datasets resolved at single-cell resolution, Visium data is captured at the level of $55$-micron-diameter spots, revealing joint expression across multiple cells. Since direct cell-type information is unavailable, we restrict evaluation to the $2$-Wasserstein distance. Moreover, since evaluation is performed midway through the $3$ time points, the IVP and next-step predictions coincide. Table 4 summarizes the comparisons between MOTFM and our methods. Consistent with our previous findings, CTF-H achieves the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
@@ 5.2 Additional Evaluations on Organogenesis and Liver Regeneration
(c) Ablations on $\lambda$
(d) OT Runtime
Figure 3: (a) and (b) compare ContextFlow with MOTFM in terms of the KL divergence between predicted and true cell type distributions under IVP sampling on two datasets, (c) summarizes the ablations on $\lambda$ in terms of the optimal choice counts, and (d) visualizes the runtime curves with respect to varying sample sizes across different OT-based coupling schemes.
5.3 Further Analysis
Cell Type Distribution. Figures 3(a) and 3(b) demonstrate the KL divergence between normalized histograms of predicted and ground-truth cell types. Compared with MOTFM, ContextFlow exhibits lower average divergence across time points, indicating that the inferred trajectories better preserve the biological composition of cell types. The cell type progression is further visualized in Figure 11 in Appendix G.4, showing the temporal trajectories of cell types evolve smoothly and consistently across consecutive developmental stages. Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development progresses, while terminal fates, including muscle, the cartilage primordium, and the liver, emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout, confirming that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
Relative Importance of Priors. We conduct extensive ablation studies on CTF-H and report them in Appendix I. To summarize trends across datasets and settings, we report the counts of the optimal $\lambda$ values in Figure 3(c). For IVP sampling, $\lambda=0.5/1$ (both useful, with a bias towards SS) are more frequently optimal settings, while for next-step sampling, $\lambda=0$ (bias towards LR) is the most frequent. This behavior can be explained by the different roles of the priors. The SS prior ($\lambda=1$) serves as a proxy for relative spatial consistency across tissue microenvironments, providing a global structural constraint that remains informative over long integration horizons and is therefore beneficial for IVP sampling. In contrast, the LR prior is informative depending on how distinct the cell-cell communication patterns are over a given time interval, making it more localized and time-dependent, particularly informative for short-range transitions, and thus more beneficial for next-step sampling.
Computational Advantage. We compare the performance of CTF-H OT, the coupling scheme used in ContextFlow, against state-of-the-art spatiotemporal OT methods, including DeST-OT (Halmos et al., 2025), MOSCOT (Klein et al., 2025), and TOAST (Ceccarelli et al., 2026). Figure 3(d) reports the runtime of different OT methods on the Brain Regeneration dataset with respect to varying sample sizes, where CTF-H OT is significantly faster than prior methods (270270x for DeST-OT, 257257x for MOSCOT, and 2424x for TOAST), while achieving comparable OT alignment performance (see Tables 6-8 in Appendix F for comparisons in growth distortion, migration, and coupled transcription distance metrics). These results suggest that the proposed prior integration strategy is more suited for flow matching frameworks, where thousands of OT problems are involved during training.